four years ago, marion was 2/3 as old as les was.now she is 3/4 as old as he is. how old is each now?
I know that i have to always compare the ages in the same time, so i have
2/3m=L
3/4m plus 4=L plus 4
I am fairly certain the second equation is wrong.

Answer by scott8148(6628)

M = "Marion NOW" ___ L = "Les NOW"

"four years ago, marion was 2/3 as old as les was" ___ M - 4 = 2/3 (L - 4)

"now she is 3/4 as old as he is" ___ M = (3/4) L

substituting ___ (3/4)L - 4 = 2/3 (L - 4)

9 L - 48 = 8L - 32 ___ L = 16

substituting ___ M = (3/4)(16) = 12